OpenStudy (anonymous):

Can someone help me simplify 2 r^3 x sqrt2 / 2 pi r^2 (r + 2) ?

OpenStudy (anonymous):

\[\pi r^2x \sqrt{2}+2 \pi r^5x \sqrt{2}\]